# GIMLET: Generalizable and Interpretable Model Learning through Embedded Thermodynamics

### v1

@@ 3.1 Overview
We demonstrate data-driven discovery of the physics model on several problems to verify our proposed GIMLET. Points of validation are to check 1) whether the free energy and/or dissipation functions are properly discovered, and 2) whether the trained physical model can be applied to other datasets. Selected problems for validations are summarized in Table˜1. For each problem, two datasets (A and B) are prepared using conventional CFD solvers. First, all the neural networks in the GIMLET are trained for the dataset A. Since the true form of the solutions is known, the trained free energy function $g\left(\phi\right)$ and/or the dissipation function $\theta\left(\boldsymbol{u}\right)$ are compared with the true function forms. To assess generalizability, the trained $g\left(\phi\right)$ and/or $\theta\left(\boldsymbol{u}\right)$ are fixed, and only the PINN is trained for dataset B.
@@ 3.2 Burgers equation
For this problem, the INN for the free energy function $g$ is not used, and the weights for the PINN and the DissipationNet are trained. For the PINN, 8 fully-connected dense hidden layers with 20 neurons each are used with $\tanh$ activation function. For the DissipationNet, 4 hidden layers with 10 neurons each are used with softplus and sigmoid activation functions. After the $1000$ epochs of the Adam optimizer stage, the SSBroyden optimizer is executed for 10 00010\,000 epochs. For the physics loss, $N_{f}=$5000$\$ of the residual points are used for the training. For better generalization, residual points are resampled in every $1000$ by the RAD method. For the data loss, the $N_{d}=$5000$\$ of data points are randomly selected from the total $N_{x}\times N_{t}=501\times 101$ generated dataset. From the remaining data points, which are not selected for data loss evaluation, $N_{e}=$1250$\$ points are selected for test evaluation. For the regularization term $\mathcal{J}_{r}$ in the loss function, $\ell_{2}$ regularization for the PINN is employed with coefficient $\beta_{\ell_{2}}=$1\text{\times}{10}^{-11}$\$. Other regularization terms in Eq.˜22d are not used and $\beta_{g}=\beta_{\theta}=\beta_{h}=0$.
@@ Results
3.3 Kuramoto-Sivashinsky equation
@@
For a PDE involving higher-order derivatives, we consider the one-dimensional Kuramoto-Sivashinsky equation expressed by
@@
$\frac{\partial\phi}{\partial t}+\alpha\phi\frac{\partial\phi}{\partial x}+\frac{\partial^{2}}{\partial{x}^{2}}\left(\frac{\delta G}{\delta\phi}\right)=0,$ (28)
@@
For this problem, the free energy functional derivative is composed by two INNs for single-valued functions as
@@ 3.4 Constitutive model for Newtonian fluids
For a multi-dimensional system, we next consider the two-dimensional steady fluid flow of a Newtonian fluid, which is governed by the following continuity and Navier-Stokes equations:
@@
$\boldsymbol{u}\cdot\nabla\boldsymbol{u}=-\nabla p+\Gamma_{\mathit{Re}}\frac{\delta\Theta}{\delta\boldsymbol{u}},$ (34a)
@@
In the practical observation of the flow field, obtaining the velocity and pressure at the same time seems to be highly difficult. For this reason, only the velocity field is used, and the pressure is not used in the following training stages.
@@
For this problem, the DissipationFunction is defined as a four-variable function as $\theta\left(\nabla\boldsymbol{u}\right)=\theta\left(u_{x},u_{y},v_{x},v_{y}\right)$, and expressed by a neural network of 4 fully-connected dense hidden layers with 10 neurons each. For the PINN, 8 fully-connected dense hidden layers with 20 neurons each are used with $\tanh$ activation function. The same structure is used for both model discovery and generalizability check.
@@ Results
Fig. 6: Results of model training for dissipation function of Newtonian fluids. The upper panels are for model discovery, and lower panels are for model application. (a,e): History of loss function, where the total value, $\mathcal{J}_{p}$ (GE), $\mathcal{J}_{d}$ (Data), test loss (Test), $\ell_{2}$ loss (L2), and loss for pressure reference (pRef) are separately plotted. (b,f): Flow field calculated by GIMLET. Streamlines (black solid lines) and velocity magnitude (color contour). (c,g): Reference flow field calculated by Nektar. (d,h): Differences of velocity magnitude between GIMLET and reference data.
@@ 3.5 Constitutive model for non-Newtonian fluids
For the viscosity function, we apply the Bird-Corss-Carreau-Yasuda model [Cross1965, Carreau1972, Yasuda1979, Bird1986, Gallagher2019]:
@@
$\boldsymbol{u}\cdot\nabla\boldsymbol{u}=-\nabla p-\frac{\delta\Theta}{\delta\boldsymbol{u}},$ (40a)
@@
Similar to the Newtonian fluid problem, the DissipationFunction is defined as a four-variable function as $\theta\left(\nabla\boldsymbol{u}\right)=\theta\left(u_{x},u_{y},v_{x},v_{y}\right)$, and expressed by a neural network of 4 fully-connected dense hidden layers with 10 neurons each. For the PINN, 8 fully-connected dense hidden layers with 20 neurons each are used with $\tanh$ activation function. The same structure is used for both model discovery and generalizability check. After the $1000$ epochs of the Adam optimizer stage, the SSBroyden optimizer is executed for 100 000100\,000 epochs with FP64 double precision. For the physics loss, $N_{f}=$30\,000$\$ of the residual points are used for the training with the RAD resampling at every 1000 epochs. For the data loss, $N_{d}=$30\,000$\$ of data points are randomly selected from the original datasets. From the remaining data points, which are not selected for data loss evaluation, $N_{e}=$7500$\$ points are selected for test evaluation. For the regularization term $\mathcal{J}_{r}$ in the loss function, $\ell_{2}$ regularization for the PINN is employed with coefficient $\beta_{\ell_{2}}=$1\text{\times}{10}^{-11}$\$. The reference point of the pressure is introduced as $\left(x_{p},y_{p}\right)=\left(0.5,1\right)$ with the weight $\beta_{h}=1$. Other regularization terms in Eq.˜22d are not used ($\beta_{g}=\beta_{\theta}=0$). To deal with the Reynolds number difference, the coefficient $\alpha_{\theta}$ is multiplied by the dissipation function term in the governing equation Eq.˜40a for the generalization check stage, and let $\beta_{\theta}$ be an additional training parameter in the optimization.
Fig. 10: Results of model training for dissipation function of non-Newtonian fluids. The upper panels are for model discovery, and lower panels are for model application. (a,e): History of loss function, where the total value, $\mathcal{J}_{p}$ (GE), $\mathcal{J}_{d}$ (Data), test loss (Test), $\ell_{2}$ loss (L2), and loss for pressure reference (pRef) are separately plotted. (b,f): Flow field calculated by GIMLET. Streamlines (black solid lines) and velocity magnitude (color contour). (c,g): Reference flow field calculated by OpenFOAM. (d,h): Differences of velocity magnitude between GIMLET and reference data.
@@ 4 Summary and Discussion
One may argue that the trained thermodynamic functions expressed by neural networks are still ‘gray-box’. After the thermodynamic functions are trained by GIMLET, the mathematical expression can be obtained by applying symbolic regressions (SRs) based on evolutionary algorithms and, as in SINDy, based on the assumption of model sparsity. Adoption of SRs for variational structure is easy to implement in comparison to the setting up SR for spatial derivatives in governing equations.
@@
Finally, although the computational experiments in this work are limited to two-dimensional settings–consistent with the available experimental apparatus–the proposed methodology is not inherently dimension-dependent. It can be extended to three-dimensional problems and, in principle, enables three-dimensional field reconstruction from two-dimensional observational data. We expect that the present framework will also pave the way for data-driven discovery of generalizable closures for complex turbulent flows.
@@ Acknowledgements
The work of S.S. has been supported by the grant-in-aid KAKENHI JP23KK0262. The work of E.K., K.S. and G.E.K. has been supported by the DOE-MMICS SEA-CROGS DE-SC0023191. The authors are grateful to Dr. Zhicheng Wang of Brown University for providing data for the flow past a cylinder. S.S. expresses appreciation for the support from M.D. Yuu Kawashima.
@@ Appendix A Hyper parameters
The hyper parameters used in the validations are summarized in Table˜2.
Table 2: The hyper parameters used in the validation cases.
@@ Appendix B Variational principle
We denote $\check{\sigma}_{ij}$ and $\boldsymbol{J}_{q}$ for the viscous stress tensor and the heat flux, respectively. And, we define the rate-of-strain tensor $e_{ij}$ as
@@
$\rho T\left(\frac{\partial s}{\partial t}+\boldsymbol{u}\cdot\nabla s\right)-\theta+\nabla\cdot\left(\boldsymbol{J}_{q}+\mu\boldsymbol{j}_{a}\right)=0,$ (57)
@@
$\begin{split}&\int_{V}dV\left\{\rho T\left(\partial_{t}s+\boldsymbol{u}\cdot\nabla\right)s-\theta+\nabla\cdot\left(\boldsymbol{J}_{q}+\mu\boldsymbol{j}_{a}\right)\right\}\\ =&\int_{V}dV\left\{\rho T\left(\partial_{t}s+\boldsymbol{u}\cdot\nabla\right)s-\check{\sigma}_{ij}\frac{\partial u_{j}}{\partial x_{i}}-\nu\left(\partial_{t}\phi+\boldsymbol{u}\cdot\nabla\right)\phi\right\}\\ =&\int_{V}dV\left\{\rho T\partial_{t}s-\nu\partial_{t}\phi-\frac{\partial x_{j}}{\partial A_{i}}\left(\rho T\frac{\partial s}{\partial x_{j}}-\nu\frac{\partial\phi}{\partial x_{j}}+\check{f}_{j}\right)\frac{\partial A_{i}}{\partial t}\right\}=0\end{split}$ (63)
@@
$\begin{split}\rho TD_{t}s&=\check{\sigma}_{ij}e_{ij}+\mu\nabla\cdot\boldsymbol{j}_{a}-\nabla\cdot\left(\boldsymbol{J}_{q}+\mu\boldsymbol{j}_{a}\right)\\ &=\check{\sigma}_{ij}e_{ij}+\mu\nabla\cdot\boldsymbol{j}_{a}-\nabla\cdot\boldsymbol{J}_{q}-\boldsymbol{j}_{a}\cdot\nabla\mu-\mu\nabla\cdot\boldsymbol{j}_{a}\\ &=\check{\sigma}_{ij}e_{ij}-\nabla\cdot\boldsymbol{J}_{q}-\boldsymbol{j}_{a}\cdot\nabla\mu=0.\end{split}$ (88)

### v2

@@ 3.1 Overview
We demonstrate data-driven discovery of the physics model on several problems to verify our proposed GIMLET. Points of validation are to check 1) whether the free energy and/or dissipation functions are properly discovered, and 2) whether the trained physical model can be transferred to other datasets. Selected problems for validations are summarized in Table˜1. For each problem, two datasets (A and B) are prepared using conventional CFD solvers. First, all the neural networks in the GIMLET are trained for the dataset A. Since the true form of the solutions is known, the trained free energy function $g\left(\phi\right)$ and/or the dissipation function $\theta\left(\boldsymbol{u}\right)$ are compared with the true function forms. To assess generalizability, the trained $g\left(\phi\right)$ and/or $\theta\left(\boldsymbol{u}\right)$ are fixed, and only the PINN is trained for dataset B.
@@ 3.2 Burgers equation
For this problem, the FreeEnergyNet is not used, and the weights for the PINN and the DissipationNet are trained. For the PINN, 8 fully-connected dense hidden layers with 20 neurons each are used with $\tanh$ activation function. For the DissipationNet, 4 hidden layers with 10 neurons each are used with softplus and sigmoid activation functions. After the $1000$ epochs of the Adam optimizer stage, the SSBroyden optimizer is executed for 10 00010\,000 epochs. For the physics loss, $N_{f}=$5000$\$ of the residual points are used for the training. For better generalization, residual points are resampled in every $1000$ by the RAD method. For the data loss, the $N_{d}=$5000$\$ of data points are randomly selected from the total $N_{x}\times N_{t}=501\times 101$ generated dataset. From the remaining data points, which are not selected for data loss evaluation, $N_{e}=$1250$\$ points are selected for test evaluation. For the regularization term $\mathcal{J}_{r}$ in the loss function, $\ell_{2}$ regularization for the PINN is employed with coefficient $\beta_{\ell_{2}}=$1\text{\times}{10}^{-11}$\$. Other regularization terms in Eq.˜22d are not used and $\beta_{g}=\beta_{\theta}=\beta_{h}=0$.
@@ Results
3.3 Kuramoto–Sivashinsky equation
@@
For a PDE involving higher-order derivatives, we consider the one-dimensional Kuramoto–Sivashinsky equation expressed by
@@
$\frac{\partial\phi}{\partial t}+\alpha\phi\frac{\partial\phi}{\partial x}-\frac{\partial^{2}}{\partial{x}^{2}}\left(\frac{\delta G}{\delta\phi}\right)=0,$ (28)
@@
For this problem, the DissipationNet is not used, and the FreeEnergyNet is composed of two INNs for single-valued functions as
@@ 3.4 Constitutive model for Newtonian fluids
For a multi-dimensional system, we next consider the two-dimensional steady fluid flow of a Newtonian fluid, which is governed by the following continuity and Navier–Stokes equations:
@@
$\boldsymbol{u}\cdot\nabla\boldsymbol{u}=-\nabla p-\Gamma_{\mathit{Re}}\frac{\delta\Theta}{\delta\boldsymbol{u}},$ (34a)
@@
Note, as described in Section˜2, the pressure calculated by the CFD is not used in the evaluation of data losses. The data loss is evaluated by only the error in the predicted and measured velocity fields.
@@
For this problem, the FreeEnergyNet is not used, and the DissipationFunction is defined as a four-variable function as $\theta\left(\nabla\boldsymbol{u}\right)=\theta\left(u_{x},u_{y},v_{x},v_{y}\right)$, and expressed by a neural network of 4 fully-connected dense hidden layers with 10 neurons each. For the PINN, 8 fully-connected dense hidden layers with 20 neurons each are used with $\tanh$ activation function. The same structure is used for both model discovery and generalizability check.
@@ Results
Fig. 6: Results of model training for dissipation function of Newtonian fluids. The upper panels are for model discovery, and the lower panels are for model application. (a,e): History of loss function, where the total value, $\mathcal{J}_{p}$ (GE), $\mathcal{J}_{d}$ (Data), test loss (Test), $\ell_{2}$ loss (L2), and loss for pressure reference (pRef) are separately plotted. (b,f): Flow field calculated by GIMLET. Streamlines (black solid lines) and velocity magnitude (color contour). (c,g): Reference flow field calculated by Nektar. (d,h): Differences of velocity magnitude between GIMLET and reference data.
@@ 3.5 Constitutive model for non-Newtonian fluids
For the viscosity function, we apply the Bird–Corss–Carreau–Yasuda model [Cross1965, Carreau1972, Yasuda1979, Bird1986, Gallagher2019]:
@@
$\boldsymbol{u}\cdot\nabla\boldsymbol{u}=-\nabla p-\Gamma_{\mathit{Re}}\frac{\delta\Theta}{\delta\boldsymbol{u}},$ (40a)
@@
Similar to the Newtonian fluid problem, the FreeEnergyNet is not used in this problem, and the DissipationFunction is defined as a four-variable function as $\theta\left(\nabla\boldsymbol{u}\right)=\theta\left(u_{x},u_{y},v_{x},v_{y}\right)$, and expressed by a neural network of 4 fully-connected dense hidden layers with 10 neurons each. For the PINN, 8 fully-connected dense hidden layers with 20 neurons each are used with $\tanh$ activation function. The same structure is used for both model discovery and generalizability check. After the $1000$ epochs of the Adam optimizer stage, the SSBroyden optimizer is executed for 100 000100\,000 epochs with FP64 double precision. For the physics loss, $N_{f}=$30\,000$\$ of the residual points are used for the training with the RAD resampling at every 1000 epochs. For the data loss, $N_{d}=$30\,000$\$ of data points are randomly selected from the original datasets. From the remaining data points, which are not selected for data loss evaluation, $N_{e}=$7500$\$ points are selected for test evaluation. For the regularization term $\mathcal{J}_{r}$ in the loss function, $\ell_{2}$ regularization for the PINN is employed with coefficient $\beta_{\ell_{2}}=$1\text{\times}{10}^{-11}$\$. The reference point of the pressure is introduced as $\left(x_{p},y_{p}\right)=\left(0.5,1\right)$ with the weight $\beta_{h}=1$. Other regularization terms in Eq.˜22d are not used ($\beta_{g}=\beta_{\theta}=0$). Similar to the Newtonian fluid problem, the coefficient $\Gamma_{\mathit{Re}}$ is multiplied by the dissipation function term in the governing equation Eq.˜40. In the discovery stage, this coefficient is fixed as $\Gamma_{\mathit{Re}}=1$. Then, for the generalization check stage, we let $\Gamma_{\theta}$ be an additional training parameter in the optimization.
Fig. 10: Results of model training for dissipation function of non-Newtonian fluids. The upper panels are for model discovery, and the lower panels are for model application. (a,e): History of loss function, where the total value, $\mathcal{J}_{p}$ (GE), $\mathcal{J}_{d}$ (Data), test loss (Test), $\ell_{2}$ loss (L2), and loss for pressure reference (pRef) are separately plotted. (b,f): Flow field calculated by GIMLET. Streamlines (black solid lines) and velocity magnitude (color contour). (c,g): Reference flow field calculated by OpenFOAM. (d,h): Differences of velocity magnitude between GIMLET and reference data.
@@ 4 Summary and Discussion
One may argue that the trained thermodynamic functions expressed by neural networks are still ‘gray-box’. After the thermodynamic functions are trained by GIMLET, the mathematical expression can be obtained by applying symbolic regressions (SRs) based on evolutionary algorithms and, as in SINDy, based on the assumption of model sparsity. Adoption of SRs for variational structure is easy to implement in comparison to setting up SR for spatial derivatives in governing equations.
@@
Finally, although the computational experiments in this work are limited to two-dimensional settings—consistent with the available experimental apparatus—the proposed methodology is not inherently dimension-dependent. It can be extended to three-dimensional problems and, in principle, enables three-dimensional field reconstruction from two-dimensional observational data. We expect that the present framework will also pave the way for data-driven discovery of generalizable closures for complex turbulent flows.
@@ Acknowledgements
The work of S.S. has been supported by the grant-in-aid KAKENHI JP23KK0262. The work of E.K., K.S., and G.E.K. has been supported by the DOE-MMICS SEA-CROGS DE-SC0023191 and, in part by AIM for Composites, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under Award No. DE-SC0023389. The authors are grateful to Dr. Zhicheng Wang of Brown University for providing data for the flow past a cylinder. S.S. expresses appreciation for the support from M.D. Yuu Kawashima.
@@ Appendix A Hyper parameters
The hyper-parameters used in the validations are summarized in Table˜2.
Table 2: The hyper-parameters used in the validation cases.
@@ Appendix B Variational principle
We denote $\check{\sigma}_{ij}$ and $\boldsymbol{q}$ for the viscous stress tensor and the heat flux, respectively. And, we define the rate-of-strain tensor $e_{ij}$ as
@@
$\rho T\left(\frac{\partial s}{\partial t}+\boldsymbol{u}\cdot\nabla s\right)-\theta+\nabla\cdot\left(\boldsymbol{q}+\mu\boldsymbol{j}_{a}\right)=0,$ (57)
@@
$\begin{split}&\int_{V}dV\left\{\rho T\left(\partial_{t}s+\boldsymbol{u}\cdot\nabla\right)s-\theta+\nabla\cdot\left(\boldsymbol{q}+\mu\boldsymbol{j}_{a}\right)\right\}\\ =&\int_{V}dV\left\{\rho T\left(\partial_{t}s+\boldsymbol{u}\cdot\nabla\right)s-\check{\sigma}_{ij}\frac{\partial u_{j}}{\partial x_{i}}-\nu\left(\partial_{t}\phi+\boldsymbol{u}\cdot\nabla\right)\phi\right\}\\ =&\int_{V}dV\left\{\rho T\partial_{t}s-\nu\partial_{t}\phi-\frac{\partial x_{j}}{\partial A_{i}}\left(\rho T\frac{\partial s}{\partial x_{j}}-\nu\frac{\partial\phi}{\partial x_{j}}+\check{f}_{j}\right)\frac{\partial A_{i}}{\partial t}\right\}=0\end{split}$ (63)
@@
$\begin{split}\rho TD_{t}s&=\check{\sigma}_{ij}e_{ij}+\mu\nabla\cdot\boldsymbol{j}_{a}-\nabla\cdot\left(\boldsymbol{q}+\mu\boldsymbol{j}_{a}\right)\\ &=\check{\sigma}_{ij}e_{ij}+\mu\nabla\cdot\boldsymbol{j}_{a}-\nabla\cdot\boldsymbol{q}-\boldsymbol{j}_{a}\cdot\nabla\mu-\mu\nabla\cdot\boldsymbol{j}_{a}\\ &=\check{\sigma}_{ij}e_{ij}-\nabla\cdot\boldsymbol{q}-\boldsymbol{j}_{a}\cdot\nabla\mu=0.\end{split}$ (88)